OpenStudy (anonymous):

Solve for x. 6sin(x/3)-3sqrt(2)=0 do I factor 3 out first? and how do I get rid of the fraction?

OpenStudy (anonymous):

add \(3\sqrt{2}\) to both sides, then divide by 6

OpenStudy (anonymous):

you get \[\sin(\frac{x}{3})=\frac{\sqrt{3}}{2}\] which should be a value you know

OpenStudy (anonymous):

first solve for \(\frac{x}{3}\) and then solve for \(x\)